Graph the following including the axis of symmetry, identify the axis of symmetry, and create a suitable table of values x on one side and y on the other.
y = x^2 - 4x

Answer by stanbon(48558)

Graph the following including the axis of symmetry, identify the axis of symmetry, and create a suitable table of values x on one side and y on the other.
y = x^2 - 4x
The axis of symmetry is x=-b/2a=-(-4)/2=2
If x=-1, y=5
If x=0, y=0
If x=1, y=-3
If x=2, y=0
If x=3, y=-3
If x=4, y=4
===============
Cheers,
Stan H.